OpenStudy (anonymous):

Consider a uniform electric field 3 x 10 ^3 i -hat N/C. (a) What is the flux of this field through a square of 10 cm on a side whose plane is parallel to yz plane? (b) What is the flux through the same square if normal to its plane makes 1 60 degree angle with x axis?

sam (.sam.):

\[\Phi=EA=(3 \times 10^3)(0.1^2)=30 Wb\]

sam (.sam.):

For b, Use \[\Phi= EAcos(\theta)\]